A car travels 65km/h and another car travels 80km/h for t+4 hours.
Find the expression in simplest form for the sum of distance.
D travelled by the two cars in terms of t
Thanks :)

Answer by algebrahouse.com(1659)


Distance = rate x time

1st car
distance = d
rate = 65
time = t
d = 65t {distance = rate x time}

2nd car
distance = d
rate = 80
time = t + 4
d = 80(t + 4) {distance = rate x time}

Distance of two cars together:
d = 65t + 80(t + 4) {added the distances}
d = 65t + 80t + 320 {used distributive property}
d = 145t + 320 {combined like terms}

For more help like this, visit: www.algebrahouse.com